OpenStudy (nesha97):

Can someone help I know m would be -3 but i have no clue what a is

OpenStudy (nesha97):

\[a^{m}\times a ^{3}=\frac{ 1 }{ a ^{2} }\]

OpenStudy (mathmale):

Have you compiled a list of "rules of exponentiation?" That might help. Can you combine a^m and a^3 into a single expression with base a? a^? Can you re-write 1 / (a^2) so that the base is still a but the exponent is negative? If you have to say "no," please review those rules of exponentiation, and then come back.

OpenStudy (sooobored):

\[a^{-1}=\frac{1}{a}\] \[a^{-2}=\frac{1}{a^2}\] \[a^{-n}=\frac{1}{a^n}\]

OpenStudy (nesha97):

im just confused because a can equal any number right?

OpenStudy (sooobored):

yea, a is a variable a could even be another variable

OpenStudy (mathmale):

Again: Please look up and review "rules of exponentiation." Decide which rule or rules applies to the situation you're dealing with here.